OpenStudy (anonymous):

Engineers designed a metal support that forms a 65 degree angel with dam find the measure of angle supplement X+65 degrees=? x+65 degrees - ? = 180 - ? X=.

OpenStudy (anonymous):

angle = 180-65

OpenStudy (anonymous):

angle = 180-65

OpenStudy (anonymous):

angle = 180-65